Circular No. 6702

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SATURN VI (TITAN) A. Marten and T. Hidayat, Observatoire de Paris-Meudon (OPM);

R. Moreno and G. Paubert, Institut de Radioastronomie Millimétrique (IRAM); B. Bezard and D. Gautier, OPM; H. Matthews, Joint Astronomy Centre, Hilo; and T. Owen, University of Hawaii, report the detection of the isotope HC^{15}N of hydrogen cyanide on Saturn VI (Titan): “The (3-2) transition of HC^{15}N centered at 258.157 GHz was observed with the IRAM 30-m radiotelescope at Granada on July 4–7. The emission line contrast recorded at a spectral resolution of 312 kHz represents a signal-to-noise ratio of 18 after 460 min of integration time on Titan. Using a second heterodyne receiver at 3-mm wavelength, the HCN(1-0) transition at 88.63 GHz was simultaneously observed during the four days of observations. The signal-to-noise ratio at the peak of the line is ∼ 48 at 1-MHz resolution. This latter measurement is in excellent agreement with the previous ones reported by Hidayat et al. (1997, Icarus 126, 170). The HC^{14}N/HC^{15}N ratio inferred from direct comparison of the observed spectra with appropriate synthetic calculations is found to be on the order of 60, within 20 percent. Assuming that this value is representative of the ^{14}N/^{15}N ratio in Titan’s atmosphere, we derive an enrichment of this ratio by a factor of ∼ 4.5 with respect to the terrestrial value of 272.”

BL LACERTAE J. Mattox, Boston University, reports: “CGRO, XTE, and ASCA ob-

servations are now either underway or planned in response to the major optical outburst of BL Lac (IAUC 6693, 6700). Observations at other wavelengths are obviously of interest. A Web page to make plans and results rapidly available to other observers during this outburst has been established

at http://bu-ast.bu.edu/~mattox/multiwavelength/bl lac.html and

will be mirrored for easier access in Asia at http://www.astro.isas.ac.

jp/~mattox/multiwavelength/bl lac.html.”

COMET C/1997 N1 (TABUR) Visual m_{1} estimates: July 9.81 UT, 10.4 (D. A. J. Seargent, The En-

trance, N.S.W., 25×100 binoculars); 11.91, 10.3 (A. Pearce, Subiaco, W. Australia, 0.2-m reflector); 14.91, 10.0 (Pearce); 16.91, 9.8 (Pearce).

1997 July 19  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green